Can anyone give me about 4 or 5 Word Problems that you use Substitution or Elimination to solve them with.
They need to be either equations or inequality word problems, you have no idea how long today i've been struggling to figure this out, i've searched read my math book and im really in a bad situation because if I dont come up with 5 i may fail my class and that wouldn't be good.. if anyone can link me or show me some problems that have what I need That would be greatly appreciated !!!

Answer by ptaylor(2198)

I WILL GIVE YOU ONE OF EACH AND YOU CAN DO THE REST. ALSO LOOK UNDER WORD PROBLEMS ON THIS SITE AND YOU SHOULD FIND LOTS.
(1). Johnny has a total of 21 apples and oranges. He has twice as many apples as oranges. How many apples and how many oranges does Johnny have?
Let x=number of oranges
And let y=number of apples
Now we are told that:
x+y=21-------------------eq1
y=2x--------------eq2
Substitute y=2x from eq2 into eq1
x+2x=21 collect like terms
3x=21 divide both sides by 3
x=7--------------------------number of oranges
from eq2, y=2x; y=2*7=14------------------number of apples
(2) The sum of two numbers is 40 and their difference is 10. Find the numbers.
Let x=one number
And let y=the other number
Now we are told that:
x+y=40--------------------------eq1
x-y=10--------------------------eq2
Add eq1 and eq2 and by doing so we eliminate y
2x=50 divide both sides by 2
x=25----------------------------------------------one number
substitute x=25 into eq1:
25+y=40 subtract 25 from both sides
25-25+y=40-25 collect like terms
y=15-------------------------------------------the other number
CK
25-15=10
25+15=40
Hope this helps--------------------------ptaylor